OpenStudy (anonymous):

Kent has two similar cylindrical pipes, Pipe A and Pipe B. The radius of Pipe A is 6 cm, and the radius of Pipe B is 2 cm. What is the ratio of the volume of Pipe A to the volume of Pipe B? (1 point) Select one: a. 3:1 b. 9:1 c. 27:1 d. 81:1 @hartnn

hartnn (hartnn):

volume is proportional to CUBE of radius so, V1/V2 = r1^3 /r2^3 = 6^3/2^3 =...

hartnn (hartnn):

\(\Large \dfrac{V_1}{V_2} = \dfrac{r_1^3}{r_2^3}= \dfrac{6^3}{2^3} = ....\)

OpenStudy (anonymous):

is it C

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

yay thnx

hartnn (hartnn):

welcome ^_^